write an equation of the line that passes through point P and parallel to the line with the given equation P(0,-1), y = -2x + 3


M=



B=



P(-7,-4),y=16


M=


B=


P(-2,1),10x+4y=-8


M=

B=


p(0,0)y=-9x-1

M=

B=





1
Expert's answer
2020-11-03T17:13:05-0500

"\displaystyle
\textsf{By the law of parallelism,}\\
\textsf{two parallel lines have}\\
\textsf{the same gradient}\\

(a)\\
P(0,-1), \, y = -2x + 3\\

m = -2\\

\frac{y - (-1)}{x - 0} = -2\\

y + 1 = -2x\\

y = -2x - 1\\


m = -2\\

B = -1\\




(b)\\P(-7,-4),\, y=16\\

m = 0\\

\frac{y - (-4)}{x - (-7)} = 0\\

y + 4 = 0\\

y = -4\\


m = 0\\

B = -4\\



(c)\\
P(-2,1),\, 10x+4y=-8\\


4y = 8 - 10x\\

y = 2 - \frac{5x}{2}\\

m = -\frac{5}{2}\\

\frac{y - 1}{x - (-2)} = -\frac{5}{2}\\

y - 1 = -\frac{5}{2}(x + 2)\\

2y - 2= -5x - 10\\

2y + 5x = -8\\

y = -4 - \frac{5x}{2}\\

m = -\frac{5}{2}\\

B = -4\\



(d)\\
p(0,0), \, y=-9x-1\\

m = -9\\

\frac{y - 0}{x - 0} = -9\\

y = -9x\\

m = -9\\

B = 0\\"